If the probability that it may rain tomorrow is 2/5, what is the probability of that it may not rain

1
Expert's answer
2022-04-05T07:30:32-0400

"P\left( no\,\,rain \right) =1-P\left( rain \right) =1-\frac{2}{5}=\frac{3}{5}"